# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 5. Pullback of algebraic divisors
In this section we fix the hermitian lattice $L$ of signature $(n-1,1)$ as in (2.5) and a unimodular hermitian $\mathcal{O}_{\boldsymbol{k}}$-lattice $\Lambda$ with signature $(m,0)$. Then the map
(5.1) $[[L]]\rightarrow[[L^{\diamond}]],\quad M\mapsto M^{\diamond}=M\obot\Lambda$
is a well-defined bijection between the genus of $L$ and the genus of $L^{\diamond}$. Indeed, the above map is injective by Witt cancellation law (we assume $D$ is odd so $2$ is not ramified in $\boldsymbol{k}$). Using strong approximation for $\mathrm{SU}(V)$, we know that there are exactly $2^{1-o(D)}h$ isometry classes in $[[L]]$ and $[[L^{\diamond}]]$. Here $o(D)$ is the number of prime divisors of $D$ and $h$ is the class number of $\boldsymbol{k}$. So the above map is also a surjection.
@@
We denote by $\mathcal{S}^{\diamond}$ the integral model of the Shimura variety determined by $L^{\diamond}$ and by $\mathcal{S}^{\diamond,*}$ its toroidal compactification as defined in Section 2.4. We also use $\mathcal{Z}^{\diamond,*}(m)$ (resp. $\mathcal{Z}^{\diamond}(m)$) to denote special divisors on $\mathcal{S}^{\diamond,*}$ (resp. $\mathcal{S}^{\diamond}$).
@@
Let $S$ be an $\mathcal{O}_{\boldsymbol{k}}$-scheme. Consider $x_{0}=(A_{0},\iota_{0},\lambda_{0})\in\mathcal{M}_{(1,0)}(S)$. Then Serre’s tensor construction carried out in [AK18, Theorem A] induces a morphism
@@
(5.3) $\varphi_{\Lambda}:\mathcal{S}\rightarrow\mathcal{S}^{\diamond},\quad(x_{0},y)\mapsto(x_{0},y\times T_{\Lambda}(x_{0})).$
Since $\mathcal{S}^{\diamond,*}$ is proper over $\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}$ ([How15, Theorem 2.5.2]) and $\mathcal{S}^{*}$ is the Zariski closure of $\mathcal{S}$, the morphism $\varphi_{\Lambda}$ extends canonically to a morphism $\mathcal{S}^{*}\rightarrow\mathcal{S}^{\diamond,*}$, denoted by the same notation.
@@
The morphism $\varphi_{\Lambda}$ constructed above is finite.
@@
It suffices to show that $\varphi_{\Lambda}$ is proper and quasi-finite. By the Néron property of abelian schemes we know $\varphi_{\Lambda}$ is proper. The quasi-finiteness of $\varphi_{\Lambda}$ comes from the finiteness of $\mathcal{M}_{(1,0)}$ over $\mathcal{O}_{\boldsymbol{k}}$. ∎
Given a proper cusp label $\Phi=(\mathfrak{m},L)\in\mathrm{Cusp}(L)$ and a hermitian $\mathcal{O}_{\boldsymbol{k}}$-lattice $\Lambda$ of signature $(m,0)$, we may produce a new cusp label
$\Phi^{\diamond}\coloneqq(\mathfrak{m},L\obot\Lambda)\in\mathrm{Cusp}(L\obot\Lambda).$
Then the image of the boundary divisor $\mathcal{S}^{*}(\Phi)$ under $\varphi_{\Lambda}$ lies in $\mathcal{S}^{\diamond,*}(\Phi^{\diamond})\subset\mathcal{S}^{\diamond,*}$. The morphism near the boundary can be described using $\mathcal{C}_{\Phi}^{*}$ as follows. Let $L_{\Phi}=\mathfrak{m}^{\bot}/\mathfrak{m}$ as in (2.10). By (2.12), we have
$\mathcal{B}_{\Phi}\cong\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}L_{\Phi},\ \mathcal{B}_{\Phi^{\diamond}}\cong\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}L_{\Phi}^{\diamond},$
where $\mathbf{A}_{0}$ is the universal elliptic curve over $\mathcal{M}_{(1,0)}$ and $\otimes$ means Serre tensor. We then have the obvious closed immersion
$\mathcal{B}_{\Phi}\rightarrow\mathcal{B}_{\Phi^{\diamond}},\quad\sum z\otimes s\mapsto\sum z\otimes(s,0),$
where $z\in\mathcal{B}_{\Phi}(S)$ for any $\mathcal{B}_{\Phi}$-scheme $S$. Recall from §2.4 that $\mathcal{C}_{\Phi^{\diamond}}^{*}$ (resp. $\mathcal{C}_{\Phi}^{*}$) is the total space of the line bundle $\mathcal{L}_{\Phi^{\diamond}}^{-1}$ (resp. $\mathcal{L}_{\Phi}^{-1}$). By their definitions, we see immediately that
$\mathcal{L}_{\Phi^{\diamond}}|_{\mathcal{B}_{\Phi}}=\mathcal{L}_{\Phi}.$
Hence we get a natural closed immersion $\mathcal{C}_{\Phi}^{*}\rightarrow\mathcal{C}_{\Phi^{\diamond}}^{*}$. The closed immersions $\mathcal{B}_{\Phi}\rightarrow\mathcal{C}_{\Phi}^{*}$ and $\mathcal{B}_{\Phi^{\diamond}}\rightarrow\mathcal{C}_{\Phi^{\diamond}}^{*}$ defined by zero sections make the following diagram Cartesian.
@@
To summarize we have the following proposition.
@@
Let $\varphi_{\Lambda}$ be the finite morphism $\mathcal{S}^{*}\rightarrow\mathcal{S}^{\diamond,*}$ defined above. It has the following properties.
@@
    $\varphi_{\Lambda}^{*}(\mathcal{S}^{\diamond,*}(\Phi^{\prime}))=\begin{cases}\mathcal{S}^{*}(\Phi)&\text{ if $\Phi^{\prime}=\Phi^{\diamond}$ for some cusp label $\Phi$ of $\mathcal{S}^{*}$},\\ 0&\text{ otherwise}.\end{cases}$
@@
Equation (5.5) is obvious from the definition of $\Omega$. Recall from §2.4 that $\mathcal{S}^{*}(\Phi)=\Delta_{\Phi}\backslash\mathcal{B}_{\Phi}$, and the completion of $\mathcal{S}^{*}$ along $\mathcal{S}^{*}(\Phi)$ is isomorphic to the completion of $\Delta_{\Phi}\backslash\mathcal{C}_{\Phi}^{*}$ along $\Delta_{\Phi}\backslash\mathcal{B}_{\Phi}$. Similar statements hold for $\mathcal{S}^{\diamond,*}$ and $\mathcal{S}^{\diamond,*}(\Phi^{\diamond})$. From (5.4) we get a Cartesian diagram of quotient stacks
[…]
By unraveling the construction of Mumford explained in [Lan13, §4.5] or [FC13, Chapter 3], the above commutative diagram gives rise to the Cartesian diagram
[…]
The proposition follows. ∎
5.1.

### v2

@@ 5.1. Morphisms betweeen unitary Shimura varieties
In this section, we fix the hermitian lattice $L$ of signature (n1,1)(n-1,1) as in (2.5), and a unimodular hermitian $\mathcal{O}_{\boldsymbol{k}}$-lattice $\Lambda$ with signature $(m,0)$. Clearly, we have a natural map
(5.1) $L]]\rightarrow[[L^{\diamond}]],\quad M\mapsto M^{\diamond}=M\obot\Lambda$
@@
We denote by $\mathcal{S}^{\diamond}$ the integral model of the Shimura variety determined by $L^{\diamond}$ and by $\mathcal{S}^{\diamond,*}$ its toroidal compactification as defined in Section 2.7. We often add the superscript $\diamond$ to already defined notations to mark objects associated to $\mathcal{S}^{\diamond}$. For example, we use $\mathcal{Z}^{\diamond,*}(m)$ (resp. $\mathcal{Z}^{\diamond}(m)$) to denote special divisors on $\mathcal{S}^{\diamond,*}$ (resp. $\mathcal{S}^{\diamond}$).
@@
Let $S$ be an $\mathcal{O}_{\boldsymbol{k}}$-scheme. Consider $(A_{0},\iota_{0},\lambda_{0},A,\iota,\lambda,\mathcal{F})\in\mathcal{S}(S)$. Then Serre’s tensor construction carried out in [AK18, Theorem A] induces a morphism
@@
(5.3) $\varphi_{\Lambda}:\mathcal{S}\rightarrow\mathcal{S}^{\diamond},(A_{0},\iota_{0},\lambda_{0},A,\iota,\lambda,\mathcal{F})\mapsto(A_{0},\iota_{0},\lambda_{0},(A,\iota,\lambda)\times T_{\Lambda}(A_{0},\iota_{0},\lambda_{0}),\mathcal{F}\times(\operatorname{Lie}A_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)).$
@@
The morphism $\varphi_{\Lambda}$ is finite and unramified.
@@
As in the proof of Proposition 4.7, the properness of $\varphi_{\Lambda}$ follows from [FC13, Chapter I, Proposition 2.7] and the valuation criterion, the unramifiedness of $\varphi_{\Lambda}$ follows from the infinitesimal criterion and Lemma 4.6. The quasi-finiteness of $\varphi_{\Lambda}|_{\mathcal{S}}$ comes from Faltings and Tate’s isogeny theorem. Indeed, $A\times T_{\Lambda}(A_{0})\cong A^{\prime}\times T_{\Lambda}(A_{0})$ implies that $A$ and $A^{\prime}$ are isogenous by looking at the Tate modules and Dieudonné modules. ∎
Lemma 5.2.
Let the notation be as above.
  1. (1)
    The map (5.1) gives a bijection between the genus of $L$ and the genus of $L^{\diamond}$.
  2. (2)
    The map $\varphi_{\Lambda}$ induces a bijection between $\pi_{0}(\mathcal{S}(\mathbb{C}))$ (the set or connected components of $\mathcal{S}(\mathbb{C})$ ) and $\pi_{0}(\mathcal{S}^{\diamond}(\mathbb{C}))$.
Proof.
Let $V=L\otimes_{\mathbb{Z}}\mathbb{Q}$, $H=U(V)$ and $H_{0}=\hbox{SU}(V)$. Then $H_{0}$ is semi-simple and simply connected, and has thus strong apparoximation property. In particular, let
$K=U(\hat{L})=\{h\in H(\mathbb{A}_{f}):\,hL=L\}$
and $K_{0}=K\cap H_{0}(\mathbb{A}_{f})$. Then $H_{0}(\mathbb{A}_{f})=H_{0}(\mathbb{A})K_{0}$. The exact sequence
$1\rightarrow H_{0}\rightarrow H\rightarrow\boldsymbol{k}^{1}\rightarrow 1$
implies that
@@
Notice that $\det K=\hat{\mathcal{O}}_{\boldsymbol{k}}^{\times}\cap\boldsymbol{k}_{f}^{1}=\hat{\mathcal{O}}_{\boldsymbol{k}}^{1}$ does not depend on $L$. So we have
[…]
So the left hand side arrow is also a bijection. Since
[…]
the above bijection proves (1).
Claim (2) follows from (1) easily. Indeed, recall from Section 2.1 that $\mathcal{S}$ is the integral model of the Shimura variety associated to
$G=\{(g_{0},g)\in\hbox{GU}(W_{0})\times\hbox{GU}(W):\,\det g_{0}=\nu(g)\}\cong\hbox{GU}(W_{0})\times U(V),$
where $W_{0}=\mathfrak{a}_{0}\otimes_{\mathbb{Z}}\mathbb{Q}$, $W=\mathfrak{a}\otimes_{\mathbb{Z}}\mathbb{Q}$, and $V=L\otimes_{\mathbb{Z}}\mathbb{Q}$ with $L=\hbox{Hom}(\mathfrak{a}_{0},\mathfrak{a})$. The isomorphism map on $U(V)$ is give by $(g_{0},g)(f)(w_{0})=gf(g_{0}^{-1}w_{0})$. So
$\pi_{0}(\mathcal{S}(\mathbb{C}))=\hbox{CL}(\boldsymbol{k})\times\pi_{0}(\hbox{Sh}(L))$
where $\hbox{Sh}(L)$ is the Shimura variety associated to $(H,K)$. It is easy to see
$\hbox{Sh}(L)(\mathbb{C})=H(\mathbb{Q})\backslash\mathcal{D}\times H(\mathbb{A}_{f})/K\cong\cup_{M\in[[L]]}\Gamma_{M}\backslash\mathcal{D}$
with
$\Gamma_{M}=\{h\in H(\mathbb{Q}):\,hM=M\}=(gKg^{-1})\cap H(\mathbb{Q})$
where $g\in H(\mathbb{A}_{f})$ and with $gL=M$. In particular, $\pi_{0}(\hbox{Sh}(L))=[[L]]$. Now (2) is clear. ∎
Remark 5.3.
The above lemma does not imply ‘cancellation law’ on unitary Hermitian lattices in general as we require the starting lattices are in the same genus in the lemma. Indeed it is not hard to find counter example to the cancellation law. Let $L_{1}$ and $L_{2}$ be two unimodular Hermitian lattices with gram matrices
$A_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\hbox{ and }A_{1}=\begin{pmatrix}1&1\\ 1&0\end{pmatrix}$
It is easy to check that $B_{1}=\hbox{Diag}(A_{1},1)$ and $B_{2}=\hbox{Diag}(A_{2},1)$ are equaivalent (i.e., $L_{1}\obot\Lambda\cong L_{2}\obot\Lambda$): $^{t}PB_{1}\bar{P}=B_{2}$ with
$P=\begin{pmatrix}0&1&1\\ 1&0&0\\ 1&0&-1\end{pmatrix}$
On the other hand, $L_{1}$ and $L_{2}$ are not (Hermitian) equivalent over $\mathbb{Z}_{2}$ if $\boldsymbol{k}/\mathbb{Q}$ is ramified at $2$, and of course are not equivalently over $\mathbb{Z}$. By the way, when $\boldsymbol{k}/Q$ is unramified at $2$, $L_{1}$ and $L_{2}$ are equivalent.
5.2. Extension of the morphism to the boundary
We first recall an extension criterion from [Lan13]. Let $X$ be a noetherian normal $\mathcal{O}_{\boldsymbol{k}}$-scheme and $Z$ is a Cartier divisor of $X$. Let $U=X\setminus Z$. Suppose $(G,\iota,\lambda,\mathcal{F})$ is a degenerating abelian scheme of signature (n1,1)(n-1,1) relative to $(X,Z,U)$ in the sense of Definition 2.6 such that there is an $\mathcal{O}_{\boldsymbol{k}}$-morphism $\varphi:U\rightarrow\mathcal{S}$ and $(A_{0},\iota_{0},\lambda_{0},G,\iota,\lambda,\mathcal{F})|_{U}$ is the pullback of the universal family over $\mathcal{S}$. Let $s$ be a geometric point of $Z$, and an $\mathcal{O}_{\boldsymbol{k}}$-morphism $\phi:\mathrm{Spec}\,V\rightarrow X$ centered at $s$ where $V$ is a complete discrete valuation ring with fraction field $K$ such that $\phi|_{\mathrm{Spec}\,K}$ factor through $U$. Then $(A_{0},\iota_{0},\lambda_{0},G,\iota,\lambda,\mathcal{F})|_{\mathrm{Spec}\,V}$ determines an element $\xi\in\mathrm{DEG}^{\Phi}_{(n-1,1)}(\mathrm{Spec}\,V,s,\mathrm{Spec}\,K)$ for a cusp label $\Phi=(\mathfrak{m}\subset M)\in\mathrm{Cusp}([[L]])$ determined by setting $\mathfrak{m}$ to be the toric part of $G_{s}$. Define $\mathfrak{n}=\mathfrak{m}^{\vee}$ and
(5.4) $\mathrm{Sym}_{\Phi}=\mathrm{Sym}^{2}_{\mathbb{Z}}(\mathfrak{n})/\langle(x\mu)\otimes\nu-\mu\otimes(\bar{x}\nu):x\in\mathcal{O}_{\boldsymbol{k}},\mu,\nu\in\mathfrak{n}\rangle.$
By Theorem 2.8, the element $\xi\in\mathrm{DEG}^{\Phi}_{(n-1,1)}(\mathrm{Spec}\,V,s,\mathrm{Spec}\,K)$ determines a degeneration data in
$\mathrm{DD}^{\Phi}_{(n-1,1)}(\mathrm{Spec}\,V,s,\mathrm{Spec}\,K),$
which in turn determines a map $B_{\Phi}:\mathrm{Sym}_{\Phi}\rightarrow K^{*}$, see [Lan13, §6.3.1.1]. Let $v$ be the canonical discrete valuation of $K$.
Lemma 5.4.
The morphism $\varphi:U\rightarrow\mathcal{S}$ extends to a (necessarily unique) morphism $\varphi:X\rightarrow\mathcal{S}^{*}$ such that $(A_{0},\iota_{0},\lambda_{0},G,\iota,\lambda,\mathcal{F})$ is the pullback of the universal family over $\mathcal{S}^{*}$ under $\varphi$ if and only if for any geometric point $s$ of $Z$ and every $\mathcal{O}_{\boldsymbol{k}}$-morphism $\phi:\mathrm{Spec}\,V\rightarrow X$ centered at $s$ as above, $v\circ B_{\Phi}$ is contained in the unique positive semi-definite cone of $\mathrm{Sym}_{\Phi}^{\vee}\otimes_{\mathbb{Z}}\mathbb{R}$.
Proof.
The lemma follows the same proof as [Lan13, Theorem 6.4.1.1(6)]. ∎
Given a proper cusp label $\Phi=(\mathfrak{m}\subset M)\in\mathrm{Cusp}([[L]])$, we may produce a new cusp label
$\Phi^{\diamond}\coloneqq(\mathfrak{m},M\obot\Lambda)\in\mathrm{Cusp}(M\obot\Lambda)\subset\mathrm{Cusp}([[L^{\diamond}]]).$
@@ 5.2. Extension of the morphism to the boundary
The morphism $\varphi_{\Lambda}:\mathcal{S}\rightarrow\mathcal{S}^{\diamond}$ extends to a morphism $\varphi_{\Lambda}:\mathcal{S}^{*}\rightarrow\mathcal{S}^{\diamond,*}$ (still denoted by the same notation) which is finite and unramified. It has the following properties.
  1. (1)
    The pullback of the universal family over $\mathcal{S}^{\diamond,*}$ under $\varphi_{\Lambda}$ is
    (5.5) $(\mathbf{A}_{0},\iota_{0},\lambda_{0},(\mathbf{G},\iota,\lambda)\times T_{\Lambda}(\mathbf{A}_{0},\iota_{0},\lambda_{0}),\mathcal{F}\times(\operatorname{Lie}\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda))$
@@
[…]
@@
To prove the proposition, we proceed in several steps.
Step 1: Construct an analogue of $\varphi_{\Lambda}$ for the formal boundary charts $\mathcal{C}^{*}_{\Phi}$. Let $L_{\Phi}=\mathfrak{m}^{\bot}/\mathfrak{m}$ be as in (2.10), then $L_{\Phi^{\diamond}}=L_{\Phi}\oplus\Lambda$. Recall from Definition 2.7 that we have the tautological degeneration data $(B,\kappa,\psi,c,c^{\vee},\tau)$ (resp. $(B^{\diamond},\kappa^{\diamond},\psi^{\diamond},c^{\diamond},c^{\diamond,\vee},\tau^{\diamond})$) over $(\mathcal{C}^{*}_{\Phi},\mathcal{B}_{\Phi},\mathcal{C}_{\Phi})$ (resp. $(\mathcal{C}^{*}_{\Phi^{\diamond}},\mathcal{B}_{\Phi^{\diamond}},\mathcal{C}_{\Phi^{\diamond}})$). Then $B^{\diamond}=B\times(A_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)$. By (2.17), we have
$\mathcal{B}_{\Phi}\cong E\otimes_{\mathcal{O}_{\boldsymbol{k}}}L_{\Phi},\ \mathcal{B}_{\Phi^{\diamond}}\cong E\otimes_{\mathcal{O}_{\boldsymbol{k}}}L_{\Phi}^{\diamond},$
where $E=\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}(\mathfrak{n},\mathbf{A}_{0})$, $\mathbf{A}_{0}$ is the universal elliptic curve over $\mathcal{M}_{(1,0)}$, and $\otimes$ means Serre tensor. We then define a closed immersion
$\mathcal{B}_{\Phi}\hookrightarrow\mathcal{B}_{\Phi^{\diamond}},\quad\sum z_{i}\otimes s_{i}\mapsto\sum z_{i}\otimes(s_{i},0),$
where $z_{i}\in E(S)$ for any $\mathcal{B}_{\Phi}$-scheme $S$. From the definition of the Poincaré bundle $P$ (resp. $P^{\diamond}$) over $B\times B^{\vee}$ (resp. $B^{\diamond}\times B^{\diamond,\vee}$), we have
(5.8) $P^{\diamond}|_{B\times B^{\vee}}=P.$
Recall from §2.7 that $\mathcal{C}_{\Phi^{\diamond}}^{*}$ (resp. $\mathcal{C}_{\Phi}^{*}$) is the total space of the line bundle $\mathcal{L}_{\Phi^{\diamond}}^{-1}$ (resp. $\mathcal{L}_{\Phi}^{-1}$). By (5.8), we have
$\mathcal{L}_{\Phi^{\diamond}}|_{\mathcal{B}_{\Phi}}=\mathcal{L}_{\Phi}.$
Hence we get a natural closed immersion
(5.9) $\varphi_{\Lambda,\Phi}:\mathcal{C}_{\Phi}^{*}\hookrightarrow\mathcal{C}_{\Phi^{\diamond}}^{*}.$
The closed immersions $\mathcal{B}_{\Phi}\hookrightarrow\mathcal{C}_{\Phi}^{*}$ and $\mathcal{B}_{\Phi^{\diamond}}\hookrightarrow\mathcal{C}_{\Phi^{\diamond}}^{*}$ defined by zero sections make the following diagram Cartesian.
(5.10) $\mathcal{B}_{\Phi}$$\mathcal{B}_{\Phi^{\diamond}}$$\mathcal{C}_{\Phi}^{*}$$\mathcal{C}_{\Phi^{\diamond}}^{*}$$\square$}
In particular we have the following equality of Cartier divisors.
(5.11) $\varphi_{\Lambda,\Phi}^{*}(\mathcal{B}_{\Phi^{\diamond}})=\mathcal{B}_{\Phi}.$
By the above construction, $c^{\diamond}$ when restricted on $\mathcal{C}^{*}_{\Phi}$ is the composition $\mathfrak{n}\xrightarrow[]{c}B\hookrightarrow B\times(\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)$ where $B\hookrightarrow B\times(\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)$ is the obvious closed immersion $z\mapsto(z,0)$. Similar statement is true for $c^{\diamond,\vee}$. Hence we have
(5.12) $((c^{\diamond,\vee}\times c^{\diamond})^{*}(\mathcal{P}^{\diamond}))|_{B\times B^{\vee}}=(c^{\vee}\times c)^{*}(\mathcal{P}).$
Step 2: Construct an analogue of $\varphi_{\Lambda}$ for the étale neighborhood $X^{(z)}$ for a geometric point $z$ of $\mathcal{C}^{*}_{\Phi}$ (hence also an étale neighborhood of $\mathcal{S}^{*}$) as in Proposition 2.9, and glue them to get the extension we want to construct. Let
(5.13) $(\mathbf{A}_{0},\iota_{0},\lambda_{0},\mathbf{G}^{(z)},\iota^{(z)},\lambda^{(z)},\mathcal{F}^{(z)})$
be the universal semi-abelian schemes over $X^{(z)}$. Consider the tuple
(5.14) $(\mathbf{A}_{0},\iota_{0},\lambda_{0},(\mathbf{G}^{(z)},\iota^{(z)},\lambda^{(z)})\times T_{\Lambda}(\mathbf{A}_{0},\iota_{0},\lambda_{0}),\mathcal{F}\times(\operatorname{Lie}\mathbf{A}_{0}\otimes_{\mathcal{O}_{\boldsymbol{k}}}\Lambda)).$
Restricting the tuple to the interior part $U^{(z)}$ of $X^{(z)}$ defines a morphism $\varphi_{\Lambda}|_{U^{(z)}}:U^{(z)}\rightarrow\mathcal{S}^{\diamond}$ which is the composition $U^{(z)}\rightarrow\mathcal{S}\xrightarrow{\varphi_{\Lambda}}\mathcal{S}^{\diamond}$.
Let $s$ be any geometric point of $X^{(z)}\setminus U^{(z)}$. Let $V$ be a complete discrete valuation ring with fraction field $K$ and $\phi:\mathrm{Spec}\,V\rightarrow X^{(z)}$ be a morphism centered at $s$ such that $\phi|_{\mathrm{Spec}\,K}$ factor through $\mathcal{S}$. By (5.4), we have
(5.15) $\mathrm{Sym}_{\Phi}=\mathrm{Sym}_{\Phi^{\diamond}}=\mathrm{Sym}^{2}_{\mathbb{Z}}(\mathfrak{n})/\langle(x\mu)\otimes\nu-\mu\otimes(\bar{x}\nu):x\in\mathcal{O}_{\boldsymbol{k}},\mu,\nu\in\mathfrak{n}\rangle.$
The pullback of the universal family (5.13) to $\mathrm{Spec}\,V$ defines a degenerating abelian scheme, hence a degeneration data (by Theorem 2.8) in
$\mathrm{DD}^{\Phi}_{(n-1,1)}(\mathrm{Spec}\,V,s,\mathrm{Spec}\,K),$
hence defines a map $B_{\Phi}:\mathrm{Sym}_{\Phi}\rightarrow K^{*}$. Let $v$ be the canonical discrete valuation of $K$, then Lemma 5.4 implies that $v\circ B_{\Phi}$ is contained in the unique positive semi-definite cone of $\mathrm{Sym}_{\Phi}^{\vee}\otimes_{\mathbb{Z}}\mathbb{R}$. Similarly the pullback of (5.14) to $\mathrm{Spec}\,V$ defines a degeneration data in
$\mathrm{DD}^{\Phi^{\diamond}}_{(n+m-1,1)}(\mathrm{Spec}\,V,s,\mathrm{Spec}\,K),$
hence also defines a map $B_{\Phi^{\diamond}}:\mathrm{Sym}_{\Phi^{\diamond}}\rightarrow K^{*}$. Unraveling the construction of $B_{\Phi}$ and $B_{\Phi^{\diamond}}$ in [Lan13] shows that the two agree under the identification (5.15). Applying Lemma 5.4 again, we know that the morphism $\varphi_{\Lambda}|_{U^{(z)}}:U^{(z)}\rightarrow\mathcal{S}^{\diamond}$ extends to a morphism $\varphi_{\Lambda}|_{X^{(z)}}:X^{(z)}\rightarrow\mathcal{S}^{\diamond,*}$ and (5.14) is the pullback of the universal family over $\mathcal{S}^{\diamond,*}$ under $\varphi_{\Lambda}|_{X^{(z)}}$. We claim that these $\varphi_{\Lambda}|_{X^{(}z)}$ together with $\varphi_{\Lambda}:\mathcal{S}\rightarrow\mathcal{S}^{\diamond}$ glue to a morphism from $\varphi_{\Lambda}:\mathcal{S}^{*}\rightarrow\mathcal{S}^{\diamond,*}$. This is true because $\mathcal{S}^{*}$ is the quotient of an étale equivalence relation (2.18) which “identifies” isomorphic semi-abelian schemes and if two tuples $\xi=(A_{0},\ldots,G,\ldots)\in X^{(z)}(S)$ and $\xi^{\prime}=(A^{\prime}_{0},\ldots,G^{\prime},\ldots)\in X^{(z^{\prime})}(S)$ are isomorphic, then so are $\varphi_{\Lambda}|_{X^{(z)}}(\xi)$ and $\varphi_{\Lambda}|_{X^{(z^{\prime})}}(\xi^{\prime})$. The extension $\varphi_{\Lambda}$ thus defined is independent of the choice of $X^{(z)}$s used in the gluing process. Obviously the we have $\varphi_{\Lambda}(\mathcal{S}^{*}(\Phi))\subset\mathcal{S}^{\diamond,*}(\Phi^{\diamond})$.
Step 3: Compare the morphisms in Step 1 and 2, and conclude the proof. Let $z$ be a geometric point of $\mathcal{B}_{\Phi}$ (also considered as a geometric point of $\mathcal{C}^{*}_{\Phi}$ and $\mathcal{S}^{*}$). Let $z^{\diamond}$ be a geometric point of $\mathcal{B}_{\Phi^{\diamond}}$ whose image under the morphism $\mathcal{B}_{\Phi^{\diamond}}\rightarrow\mathcal{S}^{\diamond,*}$ is $\varphi_{\Lambda}(z)$. Choose étale neighborhoods $X^{(z^{\diamond})}\rightarrow\mathcal{S}^{\diamond,*}$ of $z^{\diamond}$ and $X^{(z)}\rightarrow\mathcal{S}^{*}$ of $z$ respectively as in Proposition 2.9. We claim that
(5.16) $\varphi_{\Lambda,\Phi}(z)=z^{\diamond}.$
Assuming the claim and using the notations of Proposition 2.9, we can define a morphism $\varphi_{\Lambda}^{(z)}:\mathrm{Spec}\,\hat{R}_{z}\rightarrow\hat{R}_{z^{\diamond}}$ by
$\varphi_{\Lambda}^{(z)}=(\gamma^{\diamond})^{-1}\circ\varphi_{\Lambda,\Phi}|_{\mathrm{Spec}\,\hat{R}_{z}}\circ\gamma,$
where $\varphi_{\Lambda,\Phi}$ is the morphism defined in Step 1. We claim that the following diagram commutes.
(5.17) $\mathrm{Spec}\,\hat{R}_{z}$$\mathrm{Spec}\,\hat{R}_{z^{\diamond}}$$\mathcal{S}^{*}$$\mathcal{S}^{\diamond,*}.$$\scriptstyle{\varphi_{\Lambda}^{(z)}}$$\scriptstyle{i}$$\scriptstyle{j}$$\scriptstyle{\varphi_{\Lambda}}$
It follows that $\varphi_{\Lambda}$ is finite and unramified when restricted to $(\mathcal{S}^{*})^{\wedge}_{\mathcal{S}^{*}(\Phi)}$ as $\varphi_{\Lambda,\Phi}$ is.
To prove the commutativity of (5.17), let $({}^{\heartsuit}G,{}^{\heartsuit}\iota,{}^{\heartsuit}\lambda,{}^{\heartsuit}\mathcal{F})$ (resp. $({}^{\heartsuit}G^{\diamond},{}^{\heartsuit}\iota^{\diamond},{}^{\heartsuit}\lambda^{\diamond},{}^{\heartsuit}\mathcal{F}^{\diamond})$) be the Mumford family over $\mathrm{Spec}\,\hat{R}_{z}$ (resp. $\mathrm{Spec}\,\hat{R}_{z^{\diamond}}$) associated to the tautological data $(B,\kappa,\psi,c,c^{\vee},\tau)$ (resp. $(B^{\diamond},\kappa^{\diamond},\psi^{\diamond},c^{\diamond},c^{\diamond,\vee},\tau^{\diamond})$). By (5.12), we have
(5.18) $(\varphi_{\Lambda,\Phi})^{*}({}^{\heartsuit}G^{\diamond},{}^{\heartsuit}\iota^{\diamond},{}^{\heartsuit}\lambda^{\diamond},{}^{\heartsuit}\mathcal{F}^{\diamond})=({}^{\heartsuit}G,{}^{\heartsuit}\iota,{}^{\heartsuit}\lambda,{}^{\heartsuit}\mathcal{F})\times T_{\Lambda}(\mathbf{A}_{0},\iota_{0},\lambda_{0}),$
By (5.18) and Proposition 2.9, we know
$(j\circ\varphi_{\Lambda}^{(z)})^{*}(\mathbf{G}^{\diamond},\iota^{\diamond},\lambda^{\diamond},\mathcal{F}^{\diamond})=(G^{(z)},\iota^{(z)},\lambda^{(z)},\mathcal{F}^{(z)})\times T_{\Lambda}(\mathbf{A}_{0},\iota_{0},\lambda_{0}).$
where $(\mathbf{G}^{\diamond},\iota^{\diamond},\lambda^{\diamond},\mathcal{F}^{\diamond})$ is the universal object over $\mathcal{S}^{\diamond,*}$. Here we use the fact that $T_{\Lambda}(\mathbf{A}_{0},\iota_{0},\lambda_{0})$ is defined over $\mathcal{O}_{\boldsymbol{k}}$, so is invariant under $\gamma$. On the other hand, using the definition of $\varphi_{\Lambda}$, we know
$(\varphi_{\Lambda}\circ i)^{*}({}^{\heartsuit}G^{\diamond},{}^{\heartsuit}\iota^{\diamond},{}^{\heartsuit}\lambda^{\diamond},{}^{\heartsuit}\mathcal{F}^{\diamond})=(G^{(z)},\iota^{(z)},\lambda^{(z)},\mathcal{F}^{(z)})\times T_{\Lambda}(\mathbf{A}_{0},\iota_{0},\lambda_{0}).$
By the universal property of $\widehat{\mathcal{C}}^{*}_{\Phi}$ for degenerating abelian schemes in the category
$\mathrm{DEG}_{(n-1,1)}^{\Phi}(\mathrm{Spec}\,\hat{R_{z}},\mathrm{Spec}\,(\hat{R}_{z}/I_{z}),\hat{\eta}_{z})$
relative to $(\mathrm{Spec}\,\hat{R_{z}},\mathrm{Spec}\,(\hat{R}_{z}/I_{z}),\hat{\eta}_{z})$, diagram (5.17) commutes. The same argument proves claim (5.16).
Equation (5.6) is obvious from the definition of $\Omega$. By (5.11), (5.17) and the fact that $\gamma$ and $\gamma^{\diamond}$ are identities when restricted to boundaries, we have equation (5.7). The proposition is proved. ∎
5.3. Pullback of line bundles